OpenStudy (anonymous):

Find the sum of the first 8 terms of the sequence. 1, -3, -7, -11, .

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

pre cal honors

OpenStudy (mathstudent55):

First, you need to find out if this is an arithmetic or a geometric sequence. Do you know how to do that?

OpenStudy (anonymous):

arithmetic

OpenStudy (anonymous):

Do you know the rule ?

OpenStudy (mathstudent55):

Oh, you know already. Good, you're correct. In an arithmetic sequence there is a common difference. If you subtract any term from the next term, you find the common difference. Then just add the common difference to the last term to find the next one. Keep on doing that for as many terms as you need to find.

OpenStudy (anonymous):

1+(-4)=-3 -3+(-4)=-7 -7+(-4)=-11 -11+(-4)=-15 -15+(-4)=-19 -19+(-4)=-23 -23+(-4)=-27 -27+(-4)=-31

OpenStudy (anonymous):

so it's 1,-3,-7,-11,-15,-19,-23,-27